OpenStudy (anonymous):

Solve for x. 9(x – 2) = 5(2 + x)

OpenStudy (anonymous):

@kiamousekia

OpenStudy (solomonzelman):

ok, do you know how to expand parenthesis on both sides?

OpenStudy (solomonzelman):

on the left side, where it says \(9(x-2)\) you need to multiply 9 times x, and then 9 times -2, and then add the results.

OpenStudy (solomonzelman):

what do you get when you multiply 9 times x? what do you get when you multiply 9 times -2?

OpenStudy (anonymous):

Thank you @SolomonZelman

OpenStudy (solomonzelman):

Yw, but I haven't really helped you solve your problem yet

OpenStudy (solomonzelman):

please, go ahead and respond to my questions:)

OpenStudy (anonymous):

9timesx=9x 9times-2= 18

OpenStudy (anonymous):

or is it -18?

OpenStudy (solomonzelman):

yes, 9 times -2 is -18

OpenStudy (anonymous):

ok thank you :)

OpenStudy (solomonzelman):

So, this is what is going on with the left side.... \(\large\color{black}{ \displaystyle 9(x – 2) = 5(2 + x) }\) I colored the 9 in red, just to highlight it (to put emphasis on it, don't worry about that) \(\large\color{black}{ \displaystyle \color{red}{9}(x – 2) = 5(2 + x) }\) Expanding: \(\large\color{black}{ \displaystyle \color{red}{9}\times (x)~+~\color{red}{9}\times (-2)= 5(2 + x) }\) Simplifying: \(\large\color{black}{ \displaystyle 9x~+(-18)= 5(2 + x) }\) \(\large\color{black}{ \displaystyle 9x-18= 5(2 + x) }\)

OpenStudy (solomonzelman):

Now, we need to work the right side the same way \(\large\color{black}{ \displaystyle 9x-18= \color{red}{5}(2 + x) }\) \(\large\color{black}{ \displaystyle 9x-18= \color{red}{5}\times (2) ~+~\color{red}{5}\times (x) }\)

OpenStudy (solomonzelman):

please simplifiy this expression (when you are done reading my previous replies)

OpenStudy (anonymous):

5times2=10 5timesx=5x?

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle 9x-18= \color{red}{5}\times (2) ~+~\color{red}{5}\times (x) }\) and this above, becomes \(\Downarrow\) \(\large\color{black}{ \displaystyle 9x-18=10 ~+~5x }\)

OpenStudy (solomonzelman):

now, you need to isolate the x,.

OpenStudy (solomonzelman):

If you know what to do say what your next step is going to be, but if you don't know, it is alright.